What is the difference between ΔH and ΔE (on a molar basis) for the combustion of n-octane (l) at 25 degree C°?

Grade:

2 Answers

Arun
4 years ago
Dear student
 
∆H – ∆E = – 4.5 × 8.315 × 298 J = – 11.15 kJ
 
Hope it helps
In case of any difficulty, please feel free to ask.
 
Regards
Vikas TU
4 years ago
Combination of n- octane is 
C8H18 + 25/2 O2 -> 8CO2 +9H2O
del H = del E + RT deln 
del n = np - nr 
del n = 9+8 -25/2 = 4.5
del H - delE = 8.314 * 4.5 * 298 = 11.49 KJ 
Hope this helps 
Good Luck